OpenStudy (anonymous):

Consider the linear transformation T : R2 to R2 which rotates 60 degrees and refects with respect to y-axis. 1. Find the matrix A of T in standard basis of R 2

OpenStudy (anonymous):

\[A=\left[\begin{matrix}-1 & 0 \\ 0 & 1\end{matrix}\right]\left[\begin{matrix}\cos \theta & -\sin \theta \\ \sin \theta & \cos \theta\end{matrix} \right]\] The right matrix twists anticlockwise by theta radians, so just plug in pi/3. The left matrix reflects the first coordinate across the vertical axis.